OpenStudy (anonymous):

7/12y-1/4=27-5/3 what is y?

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

LOL if you reply explain please :p

OpenStudy (anonymous):

multiply both sides by 12

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Whats answer though i need to check it

OpenStudy (anonymous):

7y -3 = 27*12 -20 7y = 27*12-20-3 and you can do on your own from here

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

thanks alot!